OpenStudy (anonymous):

Change 176x^2 + 99y^2= 1584 into standard form (ellipse!)

OpenStudy (amoodarya):

divide by 1584 \[176x^2 + 99y^2= 1584 \\ \div 1584\\\frac{176x^2}{1584}+\frac{99y^2}{1584}=1\\\frac{x^2}{9}+\frac{y^2}{16}=1\]

OpenStudy (amoodarya):



OpenStudy (anonymous):

definitely need a medal for this!

OpenStudy (anonymous):

Thank you very much @amoodarya !